OpenStudy (anonymous):

2a over 3 + 1 over 5 =7 over 6 whats a?? need help asap

OpenStudy (anonymous):

a=29/20

OpenStudy (anonymous):

explain please :)

OpenStudy (anonymous):

2a/3+6/30=35/30 2a/3=29/30 20a/30=29/30 20a=29 a=1.45

OpenStudy (anonymous):

thankyouu sooo muchhh!!!!

OpenStudy (anonymous):

can you best response me :0